OpenStudy (anonymous):

Line M is represented by the following equation: x − y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? 2x − y = 26 x + y = 18 2x − 2y = 36 x − y = −28

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

lol what do you think @Daniee_Bruhh

OpenStudy (anonymous):

yep its true!

OpenStudy (anonymous):

thx

OpenStudy (daniee_bruhh):

All we need to do is plug it in. A is Correct because 36-10=26 B is incorrect because 18+10=28 C is incorrect because 36-20=16 and D is incorrect because 18-10=8 lol

OpenStudy (daniee_bruhh):

He's correct lol

OpenStudy (anonymous):

yep I just did it!!:D